OpenStudy (anonymous):

Given the parent functions f(x) = 3x + 2 and g(x) = 5x − 10, what is g(x) − f(x)? @Salmon

OpenStudy (shamil98):

Anyways. g(x) = 5x- 10 f(x) = 3x + 2 (5x-10) - (3x+2) what do you get?

OpenStudy (shamil98):

you need to distribute the -1 to (3x+ 2) first and then combine the terms

OpenStudy (anonymous):

−2(x−6)

OpenStudy (shamil98):

er, not quite. 5x - 10 - (3x+2) becomes 5x - 10 - 3x - 2 do it from here.

OpenStudy (anonymous):

ok 2x-12

OpenStudy (shamil98):

yes.

OpenStudy (shamil98):

if you were asked to simplify it further it would be 2(x-6)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then

OpenStudy (shamil98):

that's your final answer 2x - 12

OpenStudy (anonymous):

g(x) − f(x) = 5x − 8 − 3x g(x) − f(x) = 5x − 12 − 3x g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x + 12 − 5x these are the options @shamil98

OpenStudy (shamil98):

it's b.

OpenStudy (anonymous):

how

OpenStudy (shamil98):

(5x-10) - (3x+2) right? so 5x - 10 - 3x - 2 all your option did was simplify the -10-2 giving you 5x - 12 - 3x

OpenStudy (anonymous):

ohhh thank you sooooo much

OpenStudy (shamil98):

mhm